OpenStudy (anonymous):

Subtract -11 from 4.

OpenStudy (anonymous):

4-(-11)=4+11=15

OpenStudy (unklerhaukus):

15

OpenStudy (anonymous):

4+11 =15

OpenStudy (anonymous):

whenever theres 2 negatives you add so 4--11 also equals 4+11 which is 15